20:00: we are on-sky with lasers

20:30: FLAO fails, CCD39 was off at 3:29:49, we had an error in the FYSOM state machine for OPT.LOOPBUF

NOMAD 1499_0210548

  • We have error during preset for Bayside stage Z that went in timeout
  • checkRef does not move FW1 to observing position
  • SX derotator was stucked

TN AO GainSorted ascending AO CL TT Res. WF [nm rms] HO Res. WF [nm rms] AO rec DIMM ["] JCL PCL Notes
20160517_043809 0.8,0.4,0.3 T 156.428 280.842 20151212_233600 1.10873 T F
20160517_045403 1.2,0.4,0.3 T 141.508 204.045 20151212_233600 1.49505 T F
22:00: all TCS stucked, we must restart the whole ARGOS SW.

SDSS J134332.85+415503

23:30: back on sky, BCU dies again...

TN AO Gain AO CL TT Res. WF [nm rms] HO Res. WF [nm rms] AO rec DIMM ["] JCL PCL Notes
20160517_070219 0.65,0.4,0.3 T 143.144 120.357 20151212_233600 0.784412 T F
20160517_070256 0.65,0.4,0.3 T 124.591 132.018 20151212_233600 0.833998 T F
20160517_070316 0.65,0.4,0.3 T 125.736 132.649 20151212_233600 0.846622 T F
20160517_073138 0, 0, 0 T 165.472 518.73 20151212_233600 0.996642 T F
20160517_073150 0, 0, 0 T 204.074 508.621 20151212_233600 0.951001 T F
20160517_073201 0, 0, 0 T 170.87 443.735 20151212_233600 0.942786 T F
20160517_073213 0, 0, 0 T 195.972 455.793 20151212_233600 0.926961 T F
20160517_073352 1.6,0.5,0.4 T 126.093 145.97 20151212_233600 0.848632 T F
20160517_073405 1.6,0.5,0.4 T 121.04 164.731 20151212_233600 0.856134 T F
20160517_073416 1.6,0.5,0.4 T 120.527 170.32 20151212_233600 0.840184 T F
20160517_073426 1.6,0.5,0.4 T 121.858 191.782 20151212_233600 0.845074 T F
20160517_074007 1.6,0.3,0.2 T 119.576 155.734 20151212_233600 0.810332 T F
20160517_074034 1.6,0.3,0.2 T 110.792 186.339 20151212_233600 0.798665 T F
20160517_074117 1.6,0.3,0.2 T 114.506 168.411 20151212_233600 0.846453 T F
20160517_074130 1.6,0.3,0.2 T 112.644 150.547 20151212_233600 0.852999 T F
20160517_074209 1.6,0.3,0.2 T 125.444 147.755 20151212_233600 0.907313 T F
20160517_080910 1.6,0.3,0.2 T 108.31 149.406 20151212_233600 1.15622 T F
20160517_080924 1.6,0.3,0.2 T 111.239 186.187 20151212_233600 1.17337 T F
20160517_080937 1.6,0.3,0.2 T 114.429 164.97 20151212_233600 1.28588 T F
Long CL operation, no useful images taken, we spent most of the time in trying to align the MOS mask to the arc. LUCI is applying wrong offsets.

SDSS J1723+3411

2:14: we try this new pointing, but we have a lot of clouds

2:30: we must close for humidity...
Topic revision: r1 - 24 May 2016, MarcoBonaglia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback